I get questions a and be but don't understand how they got the answers for c, d and e. Please explainThe heights of adult men in a certain country are normally distributed with a mean of 69.1 inches and a standard deviation ofc. What percentage of the adult male population of this country is taller than 511? The percentage taller than 511 is 23
Answer #1

© I foot = 12 inch s, = (5x12 +11) inch inch z = 71-69.1 2.6 0.73 P(x>50) = P12 70-73) = 1-0.7673 [from Z-score table = 2